OpenStudy (anonymous):

1/2 (3x-4)< = to 3/4(x-6)+1

OpenStudy (anonymous):

Let me just make it neater.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 } (3x-4) \le \frac{ 3 }{ 4(x-6)+1 }\]

OpenStudy (anonymous):

Is it this? O_O

OpenStudy (anonymous):

3 -- (x-6)+1 4

OpenStudy (anonymous):

But the first part is right?

OpenStudy (anonymous):

yes the 1st part is right

OpenStudy (anonymous):

I gtg, sorry.

OpenStudy (anonymous):

I'll help you tomorrow.

OpenStudy (anonymous):

Just do brackets ifrst and Isolate x like the first one.

OpenStudy (anonymous):

ok